Riahhpooh:

The base of a parking charge at a mall on a Sunday is $2 per hour.If a vehicle is parked for 4 hours, only $6 is charged.What percentage is saved on the base charging for a vehicle parked for 4 hours?

TETSXPREME:

2 dollars = 1 hour 8 dollars = 4 hours she paid 6 dollars so the amount saved = 8-6 = 2 dollars percentage saved = \frac{2}{8} × 100 percentage saved = 25% According to Mr. GooGle

Jojoknph:

20% because 2 is 20% of 8 which is how much the car should have been charged if the parked there for 4 hours

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME 2 dollars = 1 hour 8 dollars = 4 hours she paid 6 dollars so the amount saved = 8-6 = 2 dollars percentage saved = \frac{2}{8} × 100 percentage saved = 25% According to Mr. GooGle \(\color{#0cbb34}{\text{End of Quote}}\) 2 dollars - 1 h 8 dollars - 4 h paid 6 dollars - 8-6=2 dollars saved 8 dollars mean 100% 2 dollars mean x% x = 2*100/8 = 25% saved good job @TETSXPREME